• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name paschke

    Log entry time 16:05:43 on October25,2002

    Entry number 89015

    keyword=Shift Summary

    We were down all shift (except for first 13 minutes) due to G0 beam setup.

    The new Q1 reset procedure has been tested. Q1 Left wouldn't come back up
    after testing, so Scott and Mark worked on it all afternoon. I believe they replaced
    the motherboard in the power supply, and it appears healthy now.

    The cryotargets have been resting all day, with low coolant consumption. The G0 target was cooled down in the morning. In preparation for this, we set the JT valves for approx 200W HPH overhead on each loop, and which was plenty to ride out the slight coolant temperature burp.